If you have a balance scale and only the four weights 1 gram (g), 3 g, 9 g, and 27 g, explain how you could show that an object had the following weights: A) 5 g B) 16 g
Hint: Weights must be added to both sides of the balance scale

Answer by Alan3354(69443)

If you have a balance scale and only the four weights 1 gram (g), 3 g, 9 g, and 27 g, explain how you could show that an object had the following weights: A) 5 g B) 16 g
-------------
For 5,
5 + 1 + 3 = 9
----------
For 16,
16 + 3 + 9 = 1 + 27